Grade 12th passPhysical Chemistry

The Charge on 1 gm ions of Al3+ is : (e is equals to 1.6 *10^-19)

Ananya vyas
8 Years agoGrade 12th pass

1 Answer

Khimraj
8 Years ago
charge on one atom of AL3+ = 3e
mole of Al3+ = 1/27 
atom of Al3+ = (1/27)*6.022*1023 = 2.23*1022
charge on these atom = 3*1.6*10-19*2.23*1022 = 1.07*104C
Hope it clears. If you like answer then please approve it.